please help!
How much water should be added to 50ml of 25% acid solution to dilute it to a 10% acid solution?
A 75 ml
B. 7.5 ml
C impossible to do
Thank you!

Answer by tutorcecilia(2152)

Use this basic formula:
(% or rate of A)(Amount of A)+ (% or rate of B)(Amount of B) = C
C can equal (% or rate of C)(A+B) or C can equal (% or rate of C)(Amount of C)
.
Change percentages to decimals: Example: 10% = .10
.
Water is 0% acid.
.
(0%)(xml) + (25%)(50ml) = (10%)(xml + 50ml)
(0)(x) + (.25)(50) = (.10)(x + 50)
0 + 12.50 = .10x + 5.0
12.50 = .10x + 5.0 [Solve for "x"]
7.5 = .10x
75ml = x
So, add 75ml of water (0% acid) to 50ml of a 25% acid to yield 125ml of a 10% acid.
.
Check by plugging in all of the variables back into the original equation:
(%A)(Amount A) + (%B)(Amount B) = (%C)(A+B)
(0%)(75) + (.25)(50) = (.10)(125)
12.5 = 12.5 [Checks out]